1. Random variables X and Y follow a joint distribution


f(x, y) = { 2,0< x<y<1  0 otherwise}


Determine the correlation coefficient between X and Y


1
Expert's answer
2021-09-07T03:02:51-0400

Let X and Y be two random variables having joint pdf is given by

"f_{X,Y}= \left\{\begin{matrix}

2,\;\;\;0\leq x\leq y\leq 1 & \\

0, \;\;\;\;\;\;\;\;otherwise &

\end{matrix}\right."

Marginal pdf of X is given by

"f_X(x) = \int^{\infty }_{-\infty }f_{X,Y}(x,y)dy \\

= \int^1_x 2dy \\

= 2(1-x)"

Marginal pdf of Y is given by

"f_Y(y)= \int^{\infty }_{-\infty }f_{X,Y}(x,y)dx \\

= \int^y_0 2dx \\

= 2y"

Mean of X is given by

"E(X) = \int^{\infty }_{-\infty }xf_X(x)dx \\

= \int^1_0 2(1-x)dx \\

E(X) = \frac{1}{3} \\

E(X^2) = \int^{\infty }_{-\infty } x^2f_X(x)dx \\

=\int^1_0 2(1-x)x^2dx \\

E(X^2) = \frac{1}{6} \\

Var(X) = E(X^2) -[E(X)]^2 \\

= \frac{1}{6} -(\frac{1}{3})^2 \\

= \frac{1}{18}"

Mean of Y is given by

"E(Y) = \int^{\infty }_{-\infty } yf_Y(y)dy \\

= \int^1_0 2y^2 dy \\

= \frac{2}{3} \\

E(Y^2) = \int^{\infty }_{-\infty } y^2 f_Y(y)dy \\

= \int^1_0 2y^3dy \\

= \frac{1}{2} \\

Var(Y) = E(Y^2) -[E(Y)]^2 \\

= \frac{1}{2} - (\frac{2}{3})^2 \\

= \frac{1}{18} \\

Cov(X,Y) =E(X,Y) -E(X)E(Y) \\

E(X,Y) = \int^1_0 \int^y_0 xyf(x,y)dxdy \\

= \int^1_0 \int^y_0 2xydxdy \\

= \int^1_0 2 y[\frac{x^2}{2}]^y_0 dy \\

= \int^1_0 y^3dy \\

=[\frac{y^4}{4}]^1_0 \\

= \frac{1}{4} \\

Cov(X,Y)=(E(X,Y) -E(X)E(Y) \\

= \frac{1}{4} - \frac{1}{3} \times \frac{2}{3} \\

= \frac{1}{36}"

The corelation coefficient is

"ρ= \frac{Cov(X,Y)}{\sqrt{Var(X)} \sqrt{Var(Y)}} \\

= \frac{1/36}{ \sqrt{1/18} \times \sqrt{1/18}} \\

= \frac{1/36}{1/18} \\

= \frac{1}{2}"